OpenStudy (anonymous):

if f(x) = (x-1)/x and (f(x+h)-f(x))/h ... what is the answer.. so lost

OpenStudy (anonymous):

\[f(x)=\frac{ x-1 }{ x }\]

OpenStudy (anonymous):

\[and \frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

lets go real real slow it is a raft of algebra the first thing you need to write is what \[f(x+h)\] is do you know it ?

OpenStudy (anonymous):

no is a fine answer, i am just asking

OpenStudy (anonymous):

i know that i need to place it exactly where x is in the problem but confused about where there is 2 functions

OpenStudy (anonymous):

@satellite73 do i put the them in like this?\[\frac{ \frac{ f(x+h)-f(x)-h }{ h } }{ \frac{ f(x+h)-f(x) }{ } }\]

OpenStudy (anonymous):

there's an h on the bottom that got cut off

OpenStudy (anonymous):

leaving me with \[\frac{ f(x+h)-f(x)-h }{ f(x+h)-f(x) }\]

OpenStudy (anonymous):

oh no no hold the phone for a second you are doing too much and too little

OpenStudy (anonymous):

i think i did it.. another way obviously.. should i get 1 as the final answer?

OpenStudy (anonymous):

like i said lets go slow first of all \(f\) is not some general \(f\) it is a specific \(f\)

OpenStudy (anonymous):

no you do not get number for an answer

OpenStudy (anonymous):

i see you are confused by this so lets go slow \[f(x)=\frac{ x-1 }{ x }\] \[f(x+h)=?\]

OpenStudy (anonymous):

you have to replace \(x\) in \(\frac{x-1}{x}\) by \(x+h\)

OpenStudy (anonymous):

making \[f(x+h)=\frac{x+h-1}{x+y}\]

OpenStudy (anonymous):

ok, understood

OpenStudy (anonymous):

damn typo \[f(x+h)=\frac{x+h-1}{x+h}\]

OpenStudy (anonymous):

now we have \[f(x+h)\] next we need \[f(x+h)-f(x)\] this is where the algebra comes in

OpenStudy (anonymous):

\[f(x+h)-f(x)=\frac{x+h-1}{x+h}-\frac{x-1}{x}\] before the algebra your next job is to subtract

OpenStudy (anonymous):

well we have a common denominator... correct?

OpenStudy (anonymous):

pellet nvm

OpenStudy (anonymous):

thats my stupid mistake

OpenStudy (anonymous):

yes of course leave everything in factored form

OpenStudy (anonymous):

i wrote x as denominator when i shouldve had x+h

OpenStudy (anonymous):

no careful here \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] the common denonminator of \(x\) and \(x+h\) is NOT \(x+h\) just like the common denominator for \(5\) and \(5+2\) is not \(7\)

OpenStudy (anonymous):

don't do any multiplying out just write it leave it in factored form

OpenStudy (anonymous):

my final answer comes out to be \[\frac{ h }{ x(x+h) }\]

OpenStudy (anonymous):

maybe let me check

OpenStudy (anonymous):

yeah good work

OpenStudy (anonymous):

so then the hs cancel and im left with 1/x+h

OpenStudy (anonymous):

now divide by \(h\) i.e. cancel the \(h\)

OpenStudy (anonymous):

what happened to the other \(x\) ?

OpenStudy (anonymous):

sorry typo

OpenStudy (anonymous):

\[\frac{ 1 }{ x(x+h) }\]

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw